I need some needed help with this.
Write an equation going thourgh the points (3,2)(-2,4)
I can only get to this point and then im stuck
4-2 / -2 - 3 = -2/5
Y=-2/5X + b
4=-2/5(-2) + B
from here i dont know what to do.

Answer by mukhopadhyay(490)

You are on the right track.
From your last step, find the value of B and it is all done.
4=-2/5(-2) + B
=> 4 = 4/5 + B
=> B = 4 - 4/5 = 16/5;
Substitute this value in the equation you formed.
Answer will be: Y=(-2/5)X + 16/5;
When simplified, 5y = -2x + 16, same as 2x+5y = 16